a rectangle is 8 feet longer than it is wide and has an area of 2900 feet squared. What are its dimensions?
Answer by checkley77(12569)

L=W+8
LW=2,900
W(W+8)=2,900
W^2+8W-2,900=0
(W-50)(W+58)=0
W-50=0
W=50 ANS. FOR THE WIDTH.
L=50+8
L=58 ANS. FOR THE LENGTH.
PROOF:
50*58=2,900
2,900=2,900